OpenStudy (anonymous):

Factor completely. If a polynomial is prime, state this. 20y^2-25y+5

OpenStudy (anonymous):

(x-1)(4x-1)

OpenStudy (anonymous):

Factor out the common term, then work on the polynomial.\[20y^2-25y+5=5(4y^2-5y+1)=5(4y-1)(y-1) \]

OpenStudy (anonymous):

i got 5*(4y^2-5y+1) 5*(4y-1)*(y-1)

OpenStudy (anonymous):

i got it thanks